OpenStudy (anonymous):

3x+2[5-4(x+1)]=6(x-3)+2(x+5)

OpenStudy (anonymous):

distribute the 4, distribute the 6, distribute the 2

OpenStudy (anonymous):

I solved and got x=16/13th's

OpenStudy (anonymous):

I got x=6/13

Directrix (directrix):

I got 10/13.

OpenStudy (anonymous):

at least we all got the same denominator! haha

OpenStudy (anonymous):

True...now to find the numerator

Directrix (directrix):

Right on. Who wants to post his or her work first for the others to trouble-shoot?

OpenStudy (anonymous):

I got 3x=2[5-4x-1]=6x-18+2x+10

OpenStudy (anonymous):

*3x+2

OpenStudy (anonymous):

then 3x+10-8x-2=6x-18+2x+10

OpenStudy (anonymous):

3x+10-8x-4=6x-18+2x+10

OpenStudy (anonymous):

ohhh -4 is where I messed up.

Directrix (directrix):

I got 3x=2[5-4x-1]= ============= 3x + 2[5 - 4x - 4] =

Directrix (directrix):

@oholmes --> And, then ... ?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

Im confused I got 2/5th's

OpenStudy (anonymous):

okay...so we have 3x+2(5-4x-4)=6x-18+2x+10. Do we agree on that part?

OpenStudy (anonymous):

yees

OpenStudy (anonymous):

lets work with the left side first...

OpenStudy (anonymous):

okay I got 3x-10-8x-8=

OpenStudy (anonymous):

3x+10-8x-8 after we distribute the 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

-5x+2

OpenStudy (anonymous):

3x-8x and 10-8

OpenStudy (anonymous):

yep. now lets work on the right side

OpenStudy (anonymous):

6x-18+2x+10

OpenStudy (anonymous):

ok, 8x-8

OpenStudy (anonymous):

-5x+2=8x-8

OpenStudy (anonymous):

subtract 8x from both sides

OpenStudy (anonymous):

i got 10/13

OpenStudy (anonymous):

-13x+2=-8; subtract 2 from from both sides

Directrix (directrix):

10/13 is the answer I got.

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

my adding/subtract was off! Sorry! :)

Directrix (directrix):

Let's work another problem together in a new thread. @oholmes Post a question if you like.